# Conditional gene genealogies given the population pedigree for a diploid Moran model with selfing

### v1

@@ 4 Main results: conditional coalescence times given the pedigree
Figure 3: The limiting conditional CDF $t\mapsto\lim_{N\to\infty}\mathbb{P}_{{\rm diff}}(N^{-2}\tau^{(N,2)}\leq t|\mathcal{A}_{N})$ for our three assumptions about $\alpha_{N}$: partial selfing, limited outcrossing, and negligible outcrossing. (Left) Partial selfing, where $\lim_{N}N(1-\alpha_{N})=\infty$. The limiting CDF is deterministic and is exactly the CDF of an exponential random variable ${\rm Exp}\left(\frac{2}{2-\alpha}\right)$; the figure takes $\alpha=1$. (Center) Five realizations of the limiting CDF for $\lambda=5$ under limited outcrossing, i.e. with $\lim_{N}N(1-\alpha_{N})=\lambda=5$. The limiting CDF is random and is described precisely in Theorem 4.3, and the corresponding survival function is what we called a “random staircase” in Table 1; cf. also Figure 1. (Right) Five realizations of the limiting CDF for the case of negligible outcrossing, where $\lim_{N}N(1-\alpha_{N})=0$. The limiting CDF is random, and it is the CDF of a random constant that is exponentially distributed with rate 2, i.e. the CDF is a Heaviside function with exponentially distributed jump time.
To describe the “random staircase” for the case of limited outcrossing with $\lambda\in(0,\infty)$ in Table 1, we first consider the lineage of a single sampled gene from a single individual $\hat{X}_{0}$. We wish to describe a process which encapsulates (i) the set of potential ancestors from whom this present-day gene may have been inherited, and (ii) the corresponding probabilities that each potential ancestor contains the gene. A realization of this process, starting from the individual $\hat{X}_{0}$ and parameterized by time-step $k\in\mathbb{Z}_{+}$ in the past, is illustrated in Figure 4.
Figure 4: A realization of the set of individuals (potential ancestors) of the sampled gene within the individual $\hat{X}_{0}$, the oval at the bottom. The three thick edges depict selfing events. The fractional number in each individual is the probability that the sampled gene lies in that individual. After $3$ time-steps, the individual to which the sample lineage belongs undergoes a fragmentation due to an outcrossing event.
Consider $n=2$ sampled genes, one from each of the individuals $\hat{X}_{0}$ and $\hat{Y}_{0}$. There is a similar description to the set of potential ancestors, but now each potential ancestor has two probabilities, one for each gene. A realization of this process $\{Z^{N}_{k}\}_{k\in\mathbb{Z}_{+}}$ is illustrated in Figure 5.
Figure 5: A realization of the process $\{Z_{k}^{N}\}_{k\in\mathbb{Z}_{+}}$, where $Z_{k}^{N}$ represents the set of potential ancestors at time-step $k$ in the past of the two genes in the sampled individuals $\hat{X}_{0}$ and $\hat{Y}_{0}$, together with the corresponding probabilities (in red and blue respectively) that each potential ancestor contains each of the sample lineages. The individuals at the bottom are $\hat{X}_{0}$ on the left and $\hat{Y}_{0}$ on the right. The initial condition is $Z_{0}=\{((1,0),1),\,((0,1),2)\}$ where the first triple $((1,0),1)$ indicates that $\hat{X}_{0}$ has probabilities/masses $(1,0)$ and label 1, and the second triple $((0,1),2)$ indicates that $\hat{Y}_{0}$ has probabilities/masses $(0,1)$ and label 2.
The discrete-time process $Z^{N}=(Z_{k}^{N})_{k\in\mathbb{Z}_{+}}$ is a fragmentation-coagulation process (Bertoin, 2006) starting with two particles $\{((1,0),1),\,((0,1),2)\}$, in which each individual (a potential ancestor) is viewed as a particle and the pairs of probabilities in the first two components within each potential ancestor are the “masses” of the particle. The third component of the particle is its label.
For each $k\in\mathbb{Z}_{+}$, the value of $Z_{k}^{N}$ is a set of 3-tuples, where the sum of all the first two components of each 3-tuple is equal to $(1,1)$ because the total probability/mass for each of the two lineages must be 1. The third component is the label of the particle, which tracks the total structure of inheritance for the process. The state space $\mathcal{P}_{m,2}$ of $Z^{N}=(Z_{k}^{N})_{k\in\mathbb{Z}_{+}}$ can be represented as
$\mathcal{P}_{m,2}:=\bigcup_{j=1}^{\infty}\mathcal{P}_{m,2}^{(j)},$ (12)
where
$\mathcal{P}_{m,2}^{(j)}:=\{(p_{i},l_{i}):1\leq i\leq j,p_{i}\in[0,1]^{2},l_{i}\in\mathbb{Z}_{+}\}.$
Here $j$ represents the number of particles, and $p_{i}$ and $l_{i}$ represent, respectively, the masses and the label of the $i$-th particle.
Lemma 4.2 says that if one unit of time is $N^{2}$ time-steps and if $N\to\infty$, then under limited outcrossing $Z^{N}$ converges weakly to a continuous-time process described as follows.
Definition 4.1 (Fragmentation-coagulation process with rate $\lambda$).
Let $\lambda\in\mathbb{R}_{+}$ be a fixed deterministic number. A fragmentation-coagulation process with rate $\lambda$ is a continuous-time process taking values in $\mathcal{P}_{m,2}$ and has right-continuous sample paths almost surely, and such that
  • (i)
    Each particle fragments independently at rate $\lambda$ into two particles with equal masses which is half of that of the fragmenting particle, and each pair of particles coagulate independently at rate $2$ into a particle whose masses are the sum of those of the two coagulating particles.
  • (ii)
    If $i$ is the label of the fragmenting particle, then the labels of the two resultant particles are $i$ and $m$, where $m$ is the smallest positive integer such that all labels are distinct; if $i$ and $j$ are the labels of the two coagulating particles, then the resultant particle has label $\min\{i,j\}$.
Let $Z_{\lambda}=(Z_{\lambda}(t))_{t\in\mathbb{Z}_{+}}$ be a fragmentation-coagulation process with rate $\lambda$, starting with two immiscible unit masses $Z_{\lambda}(0)=\{((1,0),1),\,((0,1),2)\}$; i.e. it starts with two particles with masses $(1,0)$ and $(0,1)$ and with labels 1 and 2 respectively.
The terminology in Definition 4.1 is inspired by Bertoin (2006) who discusses fragmentation and coagulation processes of unit masses. The use of the word “immiscible" reflects the fact that the coagulation of two particles with masses $(p,q)$ and $(r,s)$ into a particle with masses $(p+r,q+s)$ results in no crossover between components; i.e., the two components of the masses do not mix.
Lemma 4.2.
Suppose $\lim_{N\to\infty}N(1-\alpha_{N})=\lambda\in\mathbb{R}_{+}$. Then $(Z_{\lfloor tN^{2}\rfloor}^{N})_{t\in\mathbb{R}_{+}}$ converges in distribution in the Skorokhod space $\mathcal{D}(\mathbb{R}_{+},\mathcal{P}_{m,2})$ under $\mathbb{P}_{{\rm diff}}$ to the process $Z_{\lambda}$ in Definition 4.1.
Lemma 4.2 is proven in Section 6.1. If we ignore the masses, e.g. ignoring the pairs of probabilities in the individuals in Figure 5, then we obtain a simpler process $G^{N}=(G_{k}^{N})_{k\in\mathbb{Z}_{+}}$ from $Z^{N}$ that keeps track of the set of potential ancestors and their ancestral relationships (i.e. which splits into, and which two particles coagulate). It is the particle process consisting solely of the labels of the particles of $Z^{N}$. This process $G^{N}$ has a graphical structure that represents all possible trajectories of $(\hat{X}_{k},\hat{Y}_{k})_{k\in\mathbb{Z}_{+}}$ (recall their definition in Theorem 3.1) and can be viewed as a subset of the population pedigree $\mathcal{G}_{N}$.
Similarly, ignoring the masses in $Z_{\lambda}$, we obtain a process which we denote by $G_{\lambda}=(G_{\lambda}(t))_{t\in\mathbb{R}_{+}}$. This continuous-time process is a particle system starting with two different particles, where each particle splits into two with rate $\lambda$ and each pair of particles coagulate independently at rate $2$. The labelling process of the particles is retained from the process $Z_{\lambda}$. It is equal in distribution to an ancestral recombination graph (Griffiths, 1991; Griffiths and Marjoram, 1997) or an ancestral selection graph (Krone and Neuhauser, 1997; Neuhauser and Krone, 1997). Throughout this paper, we call $G_{\lambda}$ an ancestral graph. A realization of an ancestral graph is presented on the left in Figure 6, and further examples are offered in Figure 7.
Figure 6: (Left) A realization $G_{\lambda}$ of the random ancestral graph $G_{\lambda}$ starting with two particles/nodes. For this particular realization, the overlap times of the sample lineages are $t_{1}<t_{2}<t_{3}$. (Right) The “random staircase” mentioned in Table 1 corresponding to this realization of $G_{\lambda}$. The conditional distribution of the limiting coalescence time $T_{\lambda}$, given $G_{\lambda}$, is obtained by tracing ancestral genetic lineages backwards in time along the graph $G_{\lambda}$. Given this $G_{\lambda}$, $T_{\lambda}$ must take values in $\{t_{1},t_{2},t_{3}\}$. One can read off the conditional CDF of $T_{\lambda}$ as follows: The node of $G_{\lambda}$ at time $t_{1}$ contains $\frac{3}{4}$ of the genetics of the right lineage and the whole of that of the left lineage, so $\mathbb{P}(T_{\lambda}=t_{1}|G_{\lambda})=\frac{3}{4}$. Between $t_{1}$ and $t_{2}$, half of the left lineage splits into the left handle. At $t_{2}$, $\frac{1}{4}$ of the right lineage meets half of the left lineage, so $\mathbb{P}(T=t_{2}|G_{\lambda})=\frac{1}{8}$. The remainder of the right lineage meets the remainder of the left lineage at $t_{3}$ so $\mathbb{P}(T_{\lambda}=t_{3})=\frac{1}{8}$.
Next, we consider two conditionally independent random walks on a given ancestral graph. Given $G_{\lambda}=(G_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ and let $G_{\lambda}(0)=\{x(0),\,y(0)\}$ be the two different particles at $t=0$. We let $(x_{\lambda}(t),\,y_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ be a continuous-time process starting at $(x(0),\,y(0))$, such that
@@
    at any fragmentation event, each of $(x_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ and $(y_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ will follow each of the two paths available with equal (i.e. 1/2) probability.
In particular, $(x_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ and $(y_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ are conditionally independent random walks on the ancestral graph $G_{\lambda}$ starting at two different particles. The first meeting time of these random walks is
@@
By Theorem 4.3, $T_{\lambda}$ under the law of $\mathbb{P}(\cdot|G_{\lambda})$ is the limit law of the pairwise coalescence time of two sample lineages given the pedigree under limited outcrossing. In Figure 7 below, we illustrate two different realizations of the process $(x_{\lambda},y_{\lambda})$ on the same ancestral graph $G_{\lambda}$.
Figure 7: Two realizations of the process $(x_{\lambda},y_{\lambda})$ on the same ancestral graph $G_{\lambda}$, one on the left and the other on right. In each realization, $x_{\lambda}$ is in blue and $y_{\lambda}$ in red. At each fragmentation of the graph, where a particle splits into two, each walk chooses which particle path to follow fairly i.e. with probability $\frac{1}{2}$ each. The thick purple dot represents the first meeting time $T_{\lambda}$ of the two paths $x_{\lambda}$ and $y_{\lambda}$. For these two realizations, $T_{\lambda}=t_{1}$ on the left and $T_{\lambda}=t_{2}$ on the right, where $\{t_{1},t_{2},t_{3}\}$ are shown in Figure 6,
4.1 Rigorous statements of convergence results
We are now ready to state our main results for coalescence times in Theorems 4.3 and 4.4. Recall the sigma field $\mathcal{A}_{N}$ in (10) and the pairwise coalescence time $\tau^{(N,2)}$ defined in (8).
Theorem 4.3.
For any $t\in\mathbb{R}_{+}$, as $N\to\infty$,
$\mathbb{P}_{{\rm diff}}\left(N^{-2}\tau^{(N,2)}>t\,|\,\mathcal{A}_{N}\right)\,{\buildrel{d}\over{\longrightarrow}}\,\begin{dcases}e^{-\frac{2}{2-\alpha}t}&\quad\text{if}\quad N(1-\alpha_{N})\rightarrow\infty\;\text{ and }\,\alpha_{N}\rightarrow\alpha\in[0,1]\\ \mathbb{P}(T_{\lambda}>t\,|\,G_{\lambda})&\quad\text{if}\quad N(1-\alpha_{N})\rightarrow\lambda\in\mathbb{R}_{+}\\ \mathbbm{1}_{\{{\rm Exp}(2)>t\}}&\quad\text{if}\quad N(1-\alpha_{N})\rightarrow 0\end{dcases}.$
Theorem 4.3 offers a description of the conditional distribution of the coalescence time $\tau^{(N,2)}$ for a sample of two gene copies from different individuals in a population of size $N$ given the pedigree. Simulations of the random CDFs can be seen in Figure 3. It says that the law of $N^{-2}\tau^{(N,2)}$ for two lineages in different individuals is robust to the structure of the pedigree so long as the rate at which any individual undergoes a splitting $(1-\alpha_{N})N^{-1}$ is significantly faster than the rate at which any two sample lineages in distinct individuals coalesce, $N^{-2}+O(N^{-3})$, in other words that the ratio of the former over the latter tends to infinity. This robustness for the case of partial selfing is simply that the conditional limit agrees exactly with that in Theorem 3.1. If, on the other hand, the rate of coalescence is comparable to or dominates the rate of splitting of an individual, as in the cases of limited outcrossing and negligible outcrossing, the theorem demonstrates the non-robustness of the pairwise coalescence time to the pedigree. That is, the conditional pairwise coalescence time, more precisely its survival function or equivalently its CDF, does not converge to its mean.
From a technical point of view, Theorem 4.3 makes precise how the random probability measure $\xi_{N}:=\mathbb{P}_{{\rm diff}}\left(N^{-2}\tau^{(N,2)}\in\cdot\,|\,\mathcal{A}_{N}\right)$ converges, as $N\to\infty$, to a $\mathcal{P}(\mathbb{R}_{+})$-valued random variable $\xi$ which happens to be deterministic in the partial-selfing case and random in the other two cases. The random variable $\xi$ is the law of an ${\rm Exp}\left(\frac{2}{2-\alpha}\right)$ random variable in the first case, the conditional law of $T_{\lambda}$ given $G_{\lambda}$ in the second case, and the (random) law $\textbf{1}_{\{{\rm Exp}(2)\in\cdot\}}$ in the third case. We anticipate that, with additional effort, one could justify the above convergence in the sense of weak convergence in the space $\mathcal{P}(\mathbb{R}_{+})$ of probability measures on $\mathbb{R}_{+}$.
@@ 4.1 Rigorous statements of convergence results
$\mathbb{P}_{\rm{same}}(N^{-2}\tau^{(N,2)}>t|\mathcal{A}_{N})\,{\buildrel{d}\over{\longrightarrow}}\,\begin{dcases}e^{-2t}&\quad\text{if }\alpha=0\\ 2^{-U}e^{-2t}&\quad\text{if }\alpha\in(0,1)\\ 0&\quad\text{if }\alpha=1\end{dcases},$
@@
Remark 4.5.
As in (2), the quantity $U$ represents the number of selfing events in the ancestry of the individual from whom our two sample lineages are taken, before the first splitting or outcrossing event. It is also one less than the geometric random variable $K$ introduced in Table 1, which was the number of generations back to the first outcrossing event in ancestry of the individual.
4.2 Conditional genealogy for general sample size $n>2$
We anticipate that our main results can readily be extended to general sample size $n>2$. We do not provide details of the proofs in this paper and leave it as a conjecture in this section and for future work.
Suppose we have a sample of $n$ genes from the population at time-step 00, where $n>2$ is an arbitrary fixed number. It is customary to summarize the ancestral relationships in the sample by an ancestral process $(\Pi_{k}^{N,n})_{k\in\mathbb{Z}_{+}}$ takes values in the space $\mathcal{E}_{n}$ of partitions of $\{1,\ldots,n\}$, where indices belong to the same block of $\Pi^{N,n}_{k}$ if and only if the samples with these indices have a common ancestor $k$ steps in the past.
As before, we let $\mathbb{P}_{{\rm diff}}$ be the probability under which our $n$ samples coming from $n$ distinct individuals. Under $\mathbb{P}_{{\rm diff}}$, the initial state $\Pi_{0}^{N,n}$ of the ancestral process has $n$ blocks of singleton. We also denote by $\mathcal{A}_{N}$ the $\sigma$-algebra generated by both the pedigree and the labels of the $n$ individuals from whom we sample the $n$ sample lineages. We make the following conjecture.
@@ 4.2 Conditional genealogy for general sample size $n>2$
Under the conditional law $\mathbb{P}_{{\rm diff}}(\cdot|\mathcal{A}_{N})$, the sequence of $\mathcal{E}_{n}$-valued processes $\Pi^{N,n}=(\Pi^{N,n}_{\lfloor tN^{2}\rfloor})_{t\in\mathbb{R}_{+}}$ converges in finite dimensional distributions, as $N\to\infty$, to a random variable
@@
$\Pi=\begin{cases}\mathcal{K}^{(\alpha)}_{n}&\quad\text{if }N(1-\alpha_{N})\rightarrow\infty\quad\text{and }\alpha_{N}\rightarrow\alpha\\ P_{\lambda}\quad\text{under }\mathbb{P}(\cdot|G_{\lambda}^{n})&\quad\text{if }N(1-\alpha_{N})\rightarrow\lambda\in\mathbb{R}_{+}\end{cases}$ (14)
where $\mathcal{K}^{(\alpha)}_{n}$ is the standard Kingman $n$-coalescent with time rescaled by the factor $\frac{2}{2-\alpha}$. and $K^{(1)}_{n}$ is a single realization of $\mathcal{K}^{(1)}_{n}$.
Here we let $G_{\lambda}^{n}$ be an ancestral graph starting with $n$ initial nodes. Given $G_{\lambda}^{n}$. We let $G_{\lambda}^{n}(0)=\{x_{i}(0)\}_{1\leq i\leq n}$ be the set of the $n$ particles at $t=0$, and we let $((x_{i}(t))_{1\leq i\leq n})_{t\in\mathbb{R}_{+}}$ be the system of coalescing random walks on $G_{\lambda}^{n}$, starting at $(x_{i}(0))_{1\leq i\leq n}$ defined as before; before equation (13). Finally, $P_{\lambda}=(P_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ is a $\mathcal{E}_{n}$-valued process defined by the coalescence of the particles i.e. $i\sim_{P_{\lambda}(t)}j$ if and only if $x_{i}(t)=x_{j}(t)$.
@@
Furthermore, as in Proposition 22, we expect that as $\lambda\rightarrow\infty$, $P_{\lambda}$ under $\mathbb{P}(\cdot|G_{\lambda}^{n})$ converges to the Kingman $n$-coalescent $\mathcal{K}^{(\alpha)}_{n}$; and as $\lambda\rightarrow 0$, $P_{\lambda}$ under $\mathbb{P}(\cdot|G_{\lambda}^{n})$ converges to a single realization of $\mathcal{K}^{(1)}_{n}$, which is the limit in the negligible-outcrossing regime.
5 Properties of the scaling limit under limited outcrossing

### v2

@@ 4 Main results: conditional coalescence times given the pedigree
Figure 3: The limiting conditional CDF $t\mapsto\lim_{N\to\infty}\mathbb{P}_{{\rm diff}}(N^{-2}\tau^{(N)}\leq t|\mathcal{A}_{N})$ for our Moran model under three assumptions about $\alpha_{N}$: partial selfing, limited outcrossing, and negligible outcrossing. (Left) Partial selfing, where $\lim_{N}N(1-\alpha_{N})=\infty$. The limiting CDF is deterministic and is exactly the CDF of an exponential random variable ${\rm Exp}\left(\frac{2}{2-\alpha}\right)$, with $\alpha=1$ for comparison with those at center and right. (Center) Five realizations of the limiting CDF for $\lambda=5$ under limited outcrossing, i.e. with $\lim_{N}N(1-\alpha_{N})=\lambda=5$. The limiting CDF is random and is described precisely in Theorem 4.1. (Right) Five realizations of the limiting CDF for the case of negligible outcrossing, where $\lim_{N}N(1-\alpha_{N})=0$. The limiting CDF is random, and it is the CDF of a random constant that is exponentially distributed with rate 2, i.e. the CDF is a Heaviside function with exponentially distributed jump time.
To complete the exposition of Theorem 4.1, we now formally define the ancestral graph $G_{\lambda}$ and the pairwise coalescence time $T_{\lambda}$ of random walks theoreon.
Definition 4.3 (The ancestral graph $G_{\lambda}$).
Fix $\lambda\in\mathbb{R}_{+}$. By the ancestral graph $G_{\lambda}=\left(G_{\lambda}(t)\right)_{t\in\mathbb{R}_{+}}$ we denote a particle system beginning with two different particles, where each particle splits into two with independently with rate $\lambda$ and each pair of particles coagulate independently at rate $2$.
The ancestral graph is equal in distribution to an ancestral recombination graph (Griffiths, 1991; Griffiths and Marjoram, 1997) or an ancestral selection graph (Krone and Neuhauser, 1997; Neuhauser and Krone, 1997). Now consider coalescing random walks $x_{\lambda}$ and $y_{\lambda}$ on $G_{\lambda}$. Note that since $G_{\lambda}$ is the limiting object and $\alpha_{N}\to 1$, these walks coalesce with probability one the first time they meet. Let $G_{\lambda}(0)=\{x(0),y(0)\}$ consist of the two different particles at $t=0$. Furthermore, let $\left(x_{\lambda}(t),y_{\lambda}(t)\right)_{t\in\mathbb{R}_{+}}$ be a $G_{\lambda}$-valued continuous-time process starting at $(x(0),y(0))$ such that
@@
    at any splitting event, each of $(x_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ and $(y_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ will follow each of the two paths available with equal (i.e. 1/2) probability.
Then, $(x_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ and $(y_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ are coalescing random walks on the ancestral graph $G_{\lambda}$ starting at two different particles.
Definition 4.4 (The coalescence time $T_{\lambda}$).
The first meeting time of $x_{\lambda}$ and $y_{\lambda}$ is
@@
How one may reconstruct the CDF of the $T_{\lambda}$, conditional on a realization of $G_{\lambda},$ is discussed in Figure 4. Two realizations of $T_{\lambda}$ on the same ancestral graph are demonstrated in Figure 6.
Figure 4: (Left) A realization of the random ancestral graph $G_{\lambda}$ starting with two particles/nodes. For this particular realization, the overlap times of the sample lineages are $t_{1}<t_{2}<t_{3}$. (Right) The CDF corresponding to this $G_{\lambda}$. The conditional distribution of the limiting coalescence time $T_{\lambda}$, given $G_{\lambda}$, is obtained by tracing ancestral genetic lineages backwards in time along the graph $G_{\lambda}$. Given this $G_{\lambda}$, $T_{\lambda}$ must take values in $\{t_{1},t_{2},t_{3}\}$. One can read off the conditional CDF of $T_{\lambda}$ as follows, by tracing the ancestries of a hypothetically infinite number of unlinked loci. The node of $G_{\lambda}$ at time $t_{1}$ contains $3/4$ of the ancestries from the right sample lineage and the whole of those from the left sample lineage, so $\mathbb{P}(T_{\lambda}=t_{1}|G_{\lambda})=3/4$. Between $t_{1}$ and $t_{2}$, half of the loci following the left ancestral lineage go left at the split and the other $1/2$ continue up. The lineage on the right between $t_{1}$ and $t_{2}$ contains $1/4$ of the ancestries from the right sample lineage. So when these two lineages overlap at $t_{2}$ we have $\mathbb{P}(T_{\lambda}=t_{2}|G_{\lambda})=1/8$. Finally the remainder of the right sample lineage meets the remainder of the left sample lineage at $t_{3}$ for $\mathbb{P}(T_{\lambda}=t_{3})=1/8$.
@@
$\mathbb{P}_{\rm{same}}(N^{-2}\tau^{(N)}>t|\mathcal{A}_{N})\to\begin{dcases}e^{-t}&\quad\text{if }\alpha=0\\ 2^{-U}e^{-\frac{2}{2-\alpha}t}&\quad\text{if }\alpha\in(0,1)\\[2.25pt] 0&\quad\text{if }\alpha=1\end{dcases},$
@@
The proof of Theorem 4.5 is given in the Appendix, Section 7.2.
As in (2), the quantity $U$ represents the number of selfing events in the ancestry of the sampled individual, before the first splitting or outcrossing event. To make the connection with Theorem 4.1, the first two cases in Theorem 4.5 pertain to partial selfing (or case one) in Theorem 4.1, where is the possible for a “same” sample to enter the “diff” process. The third case in Theorem 4.5 pertains to both limited outcrossing and negligible outcrossing (or cases two and three) in Theorem 4.1, where the two gene copies in a “same” sample coalesce instantaneously with probability $1$.
4.1 Extension to larger samples
We anticipate that our main results can readily be extended to samples of arbitrary size $n>2$. We offer a detailed conjecture here but leave further exploration and proof to future work.
Suppose we have a sample of $n$ genes from the population at time-step 00, where $n>2$ is an arbitrary fixed number. As reviewed in Berestycki (2009), it is customary in formal treatments of coalescent theory to summarize the ancestral relationships of the sample by an ancestral process which we can call $(\Pi_{k}^{N,n})_{k\in\mathbb{Z}_{+}}$ and which takes values in the space $\mathcal{E}_{n}$ of partitions of $\{1,\ldots,n\}$, where indices belong to the same block of $\Pi^{N,n}_{k}$ if and only if the samples with these indices have a common ancestor $k$ steps in the past. Under $\mathbb{P}_{{\rm diff}}$, which now pertains to $n$ samples from $n$ distinct individuals, the initial state $\Pi_{0}^{N,n}$ of the ancestral process has $n$ singleton blocks. Furthermore, let $\mathcal{A}_{N}$ now be the $\sigma$-algebra generated by both the pedigree and the labels of the $n$ individuals from whom we sample the $n$ sample lineages. We make the following conjecture about diff samples.
@@ 4.1 Extension to larger samples
Under the conditional law $\mathbb{P}_{{\rm diff}}(\cdot|\mathcal{A}_{N})$, the sequence of $\mathcal{E}_{n}$-valued processes $\Pi^{N,n}=(\Pi^{N,n}_{\lfloor tN^{2}\rfloor})_{t\in\mathbb{R}_{+}}$ converges in finite dimensional distributions, as $N\to\infty$, to the random process
@@
$\Pi=\begin{cases}\mathcal{K}^{(\alpha)}_{n}&\quad\text{if }N(1-\alpha_{N})\rightarrow\infty\quad\text{and }\alpha_{N}\rightarrow\alpha\\ P_{\lambda}\quad\text{under }\mathbb{P}(\cdot|G_{\lambda}^{n})&\quad\text{if }N(1-\alpha_{N})\rightarrow\lambda\in\mathbb{R}_{+}\end{cases},$ (11)
where $\mathcal{K}^{(\alpha)}_{n}$ is the the Kingman $n$-coalescent (Kingman, 1982) with time rescaled by $\frac{2}{2-\alpha}$, $G_{\lambda}^{n}$ is an ancestral graph starting with $n$ nodes, and $P_{\lambda}=(P_{\lambda}(t))_{t\in\mathbb{R}_{+}}$ is a $\mathcal{E}_{n}$-valued process on $G_{\lambda}^{n}$ which is the extension of (10) to larger samples, so that $i\sim j$ (coalesce) if and only if $x_{i,\lambda}(t)=x_{j,\lambda}(t)$.
@@
Furthermore we expect that as $\lambda\rightarrow\infty$, $P_{\lambda}$ under $\mathbb{P}(\cdot|G_{\lambda}^{n})$ converges to the Kingman $n$-coalescent $\mathcal{K}^{(\alpha)}_{n}$ as implied by (19) below in Section 5, and that as $\lambda\rightarrow 0$, $P_{\lambda}$ under $\mathbb{P}(\cdot|G_{\lambda}^{n})$ converges to a single realization of $\mathcal{K}^{(1)}_{n}$, which is the limit in the negligible-outcrossing regime.
For the first case in (11) we further conjecture, for a sample of size $n$ containing $2m$ gene copies together as pairs in $m$ individuals and $n-2m$ gene copies each in different individuals, that each of the $m$ individuals will have its own number of selfing generations before its first outcrossing event and that these will be given by $m$ independent draws of the random variable $U$ in (2). This is a straightforward extension of the coalescent with partial selfing in Nordborg and Donnelly (1997).
5 Additional properties of limited outcrossing